Find the equation of the line tangent to the graph of f(x)=−2cos(x) at x=π4

Give your answer in point-slope form y−y0=m(x−x0). You should leave your answer in terms of exact values, not decimal approximations.
Answer #1

We are given the function,

\small f(x)=-2\cos x

Firstly, let us find the value of the function at x=π/4

\small f\left ({\pi\over 4}\right)=-2\cos {\pi\over 4 \right }=-2\times {1\over \sqrt{2}}=-\sqrt{2}

Now, we differentiating the function with respect to x, we get

\small f'(x)=-2(-\sin x)=2\sin x

Then, the value of the derivative at x=π/4 gives us the slope of the tangent at that point. So, we have

\small f'\left ( {\pi \over 4} \right )=2\sin {\pi \over 4}=2\times {1\over \sqrt {2}}=\sqrt{2}

Thus, the tangent to the graph f(x) at x=π/4 passes through the point
\small \left ( {\pi\over 4},-\sqrt{2} \right ) and has slope \small \sqrt 2

Thus the equation of the tangent, in point-slope form is

\small y-(-\sqrt{2})=\sqrt{2}\left (x-{\pi\over 4} \right )

Simplifying we get

\small y+\sqrt{2}=\sqrt{2}\left (x-{\pi\over 4} \right )

which is our required answer.